OpenStudy (anonymous):

The complex numbers z and w are represented, respectively, by the points P(x,y) and Q(u,v) respectively in Argand diagrams and w=1/z (a) show that x=u/(u^2+v^2) and find an expression for y in terms of u and v.

OpenStudy (john_es):

I would use the following, \[w=1/z\Rightarrow z=1/w\\ z=\frac{\overline{w}}{|w|^2}=\frac{u-iv}{u^2+v^2}\] \[x+iy=\frac{\overline{w}}{|w|^2}=\frac{u}{u^2+v^2}-\frac{iv}{u^2+v^2}\Rightarrow\\ \Rightarrow x=\frac{u}{u^2+v^2}\]

OpenStudy (anonymous):

zw = 1 z= x+iy w= u +vi (u+vi)(x+iy) = 1 ux +iyu +ivx -vy = 1 ux - vy =1 -----1 uy + ux = 0-----2 make y the subject in eq 2 = y = -xv/u---4 sub (y) in 4 to equation 1 and solve u get it

OpenStudy (anonymous):

for the second part make x the subject in equation 2 and substite x in equation 1 u should get \[y = \frac{-v}{v^{2}+u^{2}}\]

OpenStudy (anonymous):

Oh, got it. Thank you. @John_ES @thushananth01 ^_^

OpenStudy (anonymous):

Who do I give the medal to now? >_<

OpenStudy (anonymous):

@emcrazy14 how do you do this http://screencast.com/t/YCDyJ0fU

OpenStudy (anonymous):

What did you get for w^2?

OpenStudy (anonymous):

3 +4i

OpenStudy (anonymous):

This will be a circle with centre at (3,4) and with radius 5. You'll shade the inner region of the circle.

OpenStudy (anonymous):

 what about the other inequality